A propeller shaft has external and internal diameter 25cm and 15cm, what power can be transmitted at 110 revs/minute with a maximum shearing stress of 75 MN/m^3 and what will be twist in degrees on a 10 meter length of the shaft?.
1
Expert's answer
2020-12-22T05:07:20-0500

The polar second moment of area or the torsion constant of the cross- section about an axis through the centre


"J=\frac{\pi}{2}(r_2^4-r_1^4)=\frac{3.14}{2}(0.125^4-0.075^4)=0.334\cdot10^{-3}(m^4)"


"\tau=75(MN/m^3)"


"T=\frac{J\tau}{r_2}=\frac{0.334\cdot10^{-3}\cdot 75\cdot10^{6}}{0.125}=200400(N)"


The power generated is


"P=T\cdot\omega=200400\cdot(2\cdot3.14\cdot\frac{110}{60})=2.31\cdot10^6(W)"


The angle of twist is


"\theta=\frac{TL}{GJ}=\frac{200400\cdot10}{80\cdot10^9\cdot0.334\cdot10^{-3}}=0.075(rad)=4.3°"